OpenStudy (anonymous):

Need some help. Can someone explain the TAN function to me? I've come across it a few times. I've come to understand it but I'd like to be sure I really know what I'm saying so I don't continue to fake it till I make it

OpenStudy (kainui):

There are a bunch of ways to get to it, one way is to remember SOH CAH TOA, the TOA part means Tangent is Opposite over Adjacent, so in pictures like this:  So what it literally does is you give it an angle, and it returns to you the ratio of those two sides. Another way of looking at it is since sine and cosine both have the hypotnuse in them, you can get them to divide out that part just leaving the opposite and adjacent parts of their definition like this: \[\tan \theta = \frac{\sin \theta}{\cos \theta}\] If you have more questions feel free to ask or tag me. =)